partal derivative Kx,y) = x2 et + cosy find the fx
Answer #1

ra Au find the partial delvalve flory) = x2*4 + losy 12 fan = ? flx,y) = az eny + Losy 2 Differentiating beth sides partiallyw again, differentiating @ boll sides partially Wirto y ] alty) = x3 (28) - alsiny) s Here y is ? content BP) - x? 89 m2) + €